OpenStudy (anonymous):

i need help @goformit100 @Notamathgenius

OpenStudy (anonymous):

What is the simplified form of the expression? 5(14 - 2)2 ÷ 2 A. 60 B. 30 C. 72 D. 360

OpenStudy (goformit100):

What answer you think is correct ?

OpenStudy (notamathgenius):

PEMDAS Parenthesis Exponents Multiply Divide Add Subtract.

OpenStudy (anonymous):

ikr! @Notamathgenius but im not a math genius (/./)

OpenStudy (anonymous):

hold on...

OpenStudy (notamathgenius):

K

OpenStudy (agent0smith):

\[\Large 5(14 - 2)^2 ÷ 2\] is this your problem, or is it this \[\Large 5(14 - 2)2 ÷ 2\]

OpenStudy (anonymous):

like what do with i do with the 5? i did 14-2 and it 12 but what do i do with the 5? o.0

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (goformit100):

Yes you are correct @RoxxisaUNICORN

OpenStudy (anonymous):

help? D: @Notamathgenius

OpenStudy (anonymous):

@bakonloverk

OpenStudy (notamathgenius):

14-2 12^2 144x5 720/2=x

OpenStudy (agent0smith):

@RoxxisaUNICORN the 5 outside the brackets means multiply. \[\Large 5 \times (14 - 2)^2 ÷ 2\]

OpenStudy (anonymous):

ohhh 0.0 thanks :)

OpenStudy (goformit100):



OpenStudy (agent0smith):

\[\Large 5 \times (14 - 2)^2 ÷ 2\]first do 14-2 inside the brackets \[\Large5 \times (12)^2 ÷ 2\] Now do 12^2 \[\Large5 \times 144 ÷ 2\] Now you can do it in whatever order you like. Multiply 5 and 144, then divide by 2, or divide 144 by 2, then multiply it by 5.

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

@agent0smith

OpenStudy (texaschic101):

yes...you got it

OpenStudy (anonymous):

thnaks